OpenStudy (anonymous):

Simplify: 3x^-4w^-6z^8/4x^-2w^-20z

OpenStudy (anonymous):

@pokerstar there seems to be a mistake in what you have typed up. the expressions don't look right, especially the exponents.

OpenStudy (anonymous):

3x to the -4 power...w to the -6 power... z to the 8 power

OpenStudy (anonymous):

OH ok. So it looks like this:\[\bf \frac{ 3x^{-4} \times w^{-6} \times z^8 }{ 4x^{-2} \times w^{-20} \times z }\]Since everything is being multiplied, let's first factor out the constants and deal with the variables separately:\[\bf \frac{ 3x^{-4} \times w^{-6} \times z^8 }{ 4x^{-2} \times w^{-20} \times z }=\bf \frac{3}{4} \times \frac{ x^{-4} \times w^{-6} \times z^8 }{ x^{-2} \times w^{-20} \times z }\]Now here we make use of exponent laws. Here we will only be pretty much using the exponent law for division which says:\[\bf \frac{ a^b }{ a^c }=a^{b-c} \ | \ a \ne 0\]This just means that if the base of the numbers you are dividing is the same, then to simplify you take the base and subtract the exponents. Here is an example:\[\bf \frac{2^x}{2^n}=2^{x-n}\]Another example:\[\bf \frac{ 3^3 }{ 3^5 }=3^{3-5}=3^{-2}=\frac{ 1 }{ 9 }\]That ie exactly what you need to do in this problem. Take each letter and subtract the exponents to simplify. Can you do that? @pokerstar

OpenStudy (anonymous):

what would you get as an answer?

OpenStudy (anonymous):

\[\bf \frac{ 3 }{ 4 } x^{-4-(-2)}w^{-6-(-20)}z^{8-1}=\frac{3}{4}x^{-2}w^{14}z^7=\frac{ 3w^{14}z^7 }{ 4x^2 }\] @pokerstar

OpenStudy (anonymous):

Do you see what I did? Subtract the exponents according to the exponent law and that's the answer you get.

OpenStudy (anonymous):

Simplify: 5c^-9d^-2e^3/2c^-3d^5e^-10

OpenStudy (anonymous):

Do it the same way. You won't learn if you don't try. I told you the steps and did 1 problem for you and showed you how it's done. It's crucial for you to try and understand and solve these problems. If you however fail to, I will assist you further. @pokerstar

OpenStudy (anonymous):

for X -4 - -2 = -2 I'm confused on that?? Shouldn't ^2 be a x^-2?

OpenStudy (anonymous):

ok give me a sec and let me try this one

OpenStudy (anonymous):

I got 5e^13/2c^6d^7 is this correct?

OpenStudy (anonymous):

let me ask you this?? d^-2/d^5 how is this d^7 and not d^-7